OpenStudy (anonymous):

Find the Taylor series for f(x) centered at the given value of a. [Assume that f has a power series expansion. Do not show that R(x) → 0.] 4cos(x) at a= 3pi

OpenStudy (amistre64):

you should start by creating a list of the derivatives of f(x)

OpenStudy (anonymous):

i got the series as this: \[-4 + 0 (x-3 pi) + \frac{4}{2!}(x-3pi)^2 + 0(x-3pi)^3 - \frac{4}{2!}(x-3pi)^4+....\]

OpenStudy (anonymous):

derivatives: -4sinx -4cosx 4sinx 4cosx

OpenStudy (amistre64):

lets see if that works :) recall that 3pi is the same as pi cos(pi) = -1 ; sin(pi) = 0 cos -sin -cos sin cos -sin -cos ... we can ignore all the sin bits and the cos parts swap signs -4, 4. -4. 4. -4. 4 are the coeefs we are looking for

OpenStudy (amistre64):

lets plug those into the standard set up for the taylor: \[T_s=c_0\frac{(x-a)^0}{0!}+c_1\frac{(x-a)^1}{1!}+c_2\frac{(x-a)^2}{2!}+c_3\frac{(x-a)^3}{3!}+...\] \[T_s=-4\frac{(x-3\pi)^0}{0!}+4\frac{(x-3\pi)^1}{1!}-4\frac{(x-3\pi)^2}{2!}+4\frac{(x-3\pi)^3}{3!}+...\]

OpenStudy (amistre64):

we can simply that with summation notation

OpenStudy (amistre64):

\[T_s=\sum_{0}^{inf}\frac{(-4)^{(n+1)}(x-3pi)^n}{n!}\]

OpenStudy (anonymous):

ahh ok. Thank you!

OpenStudy (amistre64):

i see an error

OpenStudy (amistre64):

i forgot to modify my exponents when I deleted the sins

OpenStudy (phi):

I was going to say, you want even powers of n...

OpenStudy (amistre64):

make the the bottom 2n! and that x^(2n)

OpenStudy (amistre64):

:) good catch phi

OpenStudy (phi):

tyler looks good except for a typo on the last term 4! (not 2!)

OpenStudy (anonymous):

Ohh ok, Thank you both!!

OpenStudy (phi):

Getting the correct sign seems tricky

OpenStudy (anonymous):

Yeah in my first answer i had (-1)^n and not n+1. Thanks for the help

OpenStudy (phi):

tweak amistre's answer, the 4 does not belong inside the exponent (n+1)